Write a matrix formula for the perspective projection of the point (?,?,?) to the plane ?+2?+3?=0 with center of projection (−5,0,0).

This is all the info it gives.
Answer #1

We have to find the formula for the perspective projection of the point (2,4, 2) to the plane & + 2y +32 = 0 with centre of p(0.4.2) to the plane X+ 24 + 32 : 0 le 1=-24-32 plane sfp PC4.4.2) - P [24 - 32,4.2) using the homogeneous Co-ordinaty o ovo